# Unlocking Continual Learning Abilities in Language Models

### v1

@@ Abstract
Wenyu Du1thanks:   Equal Contributions.thanks:   Work done during interning at NVIDIA.  Shuang Cheng2∗  Tongxu Luo3  Zihan Qiu4  Zeyu Huang5 Ka Chun Cheung6  Reynold Cheng1thanks:   Corresponding Authors.  Jie Fu733footnotemark: 3 1The University of Hong Kong  2ICT, Chinese Academy of Sciences  3CUHK-SZ 4Tsinghua University  5University of Edinburgh  6NVIDIA  7HKUST wydu@cs.hku.hk  chengshuang22@mails.ucas.ac.cn
@@ 4 Experiments
CL Benchmark Standard Long
@@
Table 2: Average accuracy on standard CL benchmark (Order 1,2,3) and long CL benchmark (Order 4,5,6) with T5-large model. The top block contains CL methods with extra old task data or task labels, while the bottom does not.
@@ 4.1 Continual Finetuning on T5-large
We separate the baselines into two categories: without old data or task information and with old data or task information during training. For the first category, we include vanilla FT, which trains all model parameters on a sequence of tasks, and vanilla LoRA, in which fixed-size LoRA parameters are trained on a sequence of tasks. For the second category, we have rehearsal-based approaches: LoRAReplay that trains new tasks on LoRA with mixing a 2% past task, LFPT5 Qin and Joty (2021) continuously trains a soft prompt that simultaneously learns to solve the tasks and generate training samples for experience replay; architecture-based approaches: IncLoRA that incremental learning of new LoRA parameters on a sequential series of tasks, MoELora Luo et al. (2024), a vanilla MoE with LoRA number equals to the task number, SAPT-LoRA Zhao et al. (2024) extends IncLoRA by aligning learning process and selection process of LoRA, and MoCL Wang et al. (2024c) continually adds new modules and composes them with existing modules; parameter-based approaches OIncLoRA Wang et al. (2023b) [^3] O-LoRA is original name, we rename it to OIncLoRA to emphasize it is build upon IncLoRA and align with our notation. extends IncLoRA to learn different LoRAs into orthogonal subspaces.
@@ 4.2 Continual Pre-training on RoBERTa
In contrast to the previous continual finetuning setting, Ke et al. (2023) introduces DAS, a new benchmark for continual pre-training of LMs. DAS is composed of six unlabeled domain corpora, which contain three review domains and three academic paper domains. It is then evaluated using six corresponding classification datasets. Please refer to the Appendix B.1 for the details.
@@
For continual pre-training, we utilize MF1 (Macro-F1) and ACC (Accuracy) following Ke et al. (2023) to evaluate the performance after pre-training on the last domain.
@@
Table 3: Average MF1, ACC on the DAS benchmark after continual pre-training on all domains and finetuning on their corresponding end-task datasets. The top block contains CL methods with extra old task data or task labels, while the bottom does not. DAS is the average result we reproduced in 3 runs.
@@ 5 Discussions
We then provide ablations on gradient mask threshold and components as well as a visualization.
@@ 5.1 Ablation on Gradient Mask Threshold
We plot all five curves of our approach (+MIGU) for gradient mask threshold from 0.0 to 0.9 in Section 4.1. The optimal threshold value for FT+MIGU, LoRA+MIGU, and IncLoRA+MIGU settings is 0.7 while LoRAReplay+MIGU is 0.4 as shown in Figure 5444The ablation on threshold search only reports one run, so it does not align with the results in Section 4.1.. OIncLoRA+MIGU is only 0.1, which may due to the parameter updating regularized by the OIncLoRA method itself. The optimal value for IncLoRA+MIGU is 0.6, close to FT+MIGU, LoRA+MIGU, and IncLoRA+MIGU settings.
Surprisingly, with only 5% ($T=0.95$) or 1% ($T=0.99$) parameters updating, LoRA+MIGU still beats LoRA by a wide margin. This interesting finding may indicate that only a small proportion of proportional weights with large magnitudes is crucial for successful CL settings, which may be worth future investigation.
@@ 7 Limitations
We acknowledge two limitations for this work. Due to computation limitations, although we finetune Llama2-7B with LoRA, we are unable to scale our experiments to LM continual pre-training or full tuning. However, our experimental performance on continual pre-training RoBERTa indicates the great potential for the scalability of this general approach. Another limitation is we only explore an approach for unlocking the inherent CL potential of LMs through updating the gradient by the magnitude of output. There exists more discussions on exploiting innate features such as activation sparsity as discussed in the Related Work section. These limitations can be further addressed in future work.
@@ B.1 Dataset Information
Standard CL Benchmark (4 tasks) Longer CL Benchmark (15 tasks)
@@
Table 8: Summary of the results on two standard CL benchmarks with T5-large model. Averaged accuracy after training on the last task is reported. All results in the last block are averaged over 3 runs. (We reuse the table template and experiment results from O-LoRA Wang et al. (2023b) to construct the results of the top two blocks). It is noticeable some baselines in some previous literature show significant lower performance than ours, we assume this may due to our strict parameter grid search for baseline methods.
@@ D.4 Visualization
To investigate how our method enhances model performance, we visualized the variation in product magnitudes between an FT model and an FT model augmented with our MIGU technique in Figures 23,31. We employed heatmaps to depict the similarity in product magnitude distributions across different tasks. Our findings reveal that task similarity in the FT model with MIGU implementation is markedly reduced. This suggests that the models trained with our method exhibit more distinctive weight activations for different tasks, thereby mitigating their conflict. This distinction in activation patterns indicates our method’s ability to foster more task-specific representations within the model, contributing to its improved performance across varied learning scenarios.

### v2

@@ Abstract
Wenyu Du1thanks:   Equal Contributions.thanks:   Work done during interning at NVIDIA.  Shuang Cheng2∗  Tongxu Luo3  Zihan Qiu4  Zeyu Huang5 Ka Chun Cheung6  Reynold Cheng1thanks:   Corresponding Authors.  Jie Fu733footnotemark: 3 1School of Computing and Data Science, The University of Hong Kong 2ICT, Chinese Academy of Sciences  3CUHK-SZ 4Tsinghua University  5University of Edinburgh  6NVIDIA  7HKUST wydu@cs.hku.hk  chengshuang22@mails.ucas.ac.cn
@@ 4 Experiments
Methods Standard Long
@@
Table 2: Average accuracy on standard CL benchmark (Order 1,2,3) and long CL benchmark (Order 4,5,6) with T5-large model. The top block contains CL methods with extra old task data or task labels, while the bottom does not. Methods denoted with are copied from previous papers Wang et al. (2023b, 2024c); Zhao et al. (2024).
@@ 4.1 Continual Finetuning on T5-large
We separate the baselines into two categories: without old data or task information and with old data or task information during training. For the first category, we include vanilla FT, which trains all model parameters on a sequence of tasks, and vanilla LoRA, in which fixed-size LoRA parameters are trained on a sequence of tasks. For the second category, we have rehearsal-based approaches: LoRAReplay that trains new tasks on LoRA with mixing a 2% past task, LFPT5 Qin and Joty (2021) continuously trains a soft prompt that simultaneously learns to solve the tasks and generate training samples for experience replay; architecture-based approaches: IncLoRA that incremental learning of new LoRA parameters on a sequential series of tasks, MoELora Luo et al. (2024), a vanilla MoE with LoRA number equals to the task number, SAPT-LoRA Zhao et al. (2024) extends IncLoRA by aligning learning process and selection process of LoRA, and MoCL Wang et al. (2024c) continually adds new modules and composes them with existing modules; parameter-based approaches OIncLoRA Wang et al. (2023b) [^3] O-LoRA is original name, we rename it to OIncLoRA to emphasize it is build upon IncLoRA and align with our notation. extends IncLoRA to learn different LoRAs into orthogonal subspaces. Moreover, we have one multi-task learning baseline MTL from Wang et al. (2023b) as the referenced “upper bound” for the benchmark.
@@ 4.2 Continual Pre-training on RoBERTa
In contrast to the previous continual finetuning setting, Ke et al. (2023) introduces DAS, a new benchmark for continual pre-training (CPT) of LMs. DAS is composed of six unlabeled domain corpora, which contain three review domains and three academic paper domains. It is then evaluated using six corresponding classification datasets. Unlike continual finetuning, CPT is carried out in two stages: 1) continual sequential pre-training on each domain, and 2) separately continual finetuning for end tasks in each domain. [^4] Let’s take DAS as an example. Stage 1: sequentially continual pre-train Roberta on domains 1-6; stage 2: duplicate Roberta after stage 1 into six copies and continual finetune each copy on separate domain-relevant task from 1 to 6 respectively. Please refer to the Appendix B.1 for the details.
@@
For continual pre-training, we utilize MF1 (Macro-F1) and ACC (Accuracy) following Ke et al. (2023) to evaluate the performance after pre-training on the last domain. For the details, please refer to Ke et al. (2023).
@@
Table 3: Average MF1, ACC on the DAS benchmark after continual pre-training on all domains and finetuning on their corresponding end-task datasets. The top block contains CL methods with extra old task data or task labels, while the bottom does not. Methods denoted with are copied from original papers.
@@ 5 Discussions
We then provide ablations on gradient mask threshold and components as well as a visualization. We also ablate the various methods for finding important weights in Table 11 of Appendix.
@@ 5.1 Ablation on Gradient Mask Threshold
Because the current instruction finetuning datasets like standard CL benchmarks, typically lack a separate development set, we perform an ablation study to manually create a development set of 1,000 samples for each task from the training data in the standard CL benchmark. We use this new split dataset to evaluate the mask threshold selections for the FT, FT+MIGU, LoRA, and LoRA+MIGU methods. Our findings are presented in Table 5, indicating that a threshold of 0.7 is optimal for both FT+MIGU and LoRA+MIGU. Additional experiments on mask thresholds for various methods can be found in Appendix F.
Figure 5: Ablation study on the gradient mask threshold. The curves illustrate that the optimal value is concentrated around 0.7 for FT+MIGU and LoRA+MIGU.
@@ 7 Limitations
We acknowledge two limitations for this work. Due to computation limitations, although we finetune Llama2-7B with LoRA, we are unable to scale our experiments to LM continual pre-training or full tuning. However, our experimental performance on continual pre-training RoBERTa indicates the great potential for the scalability of this general approach. Another limitation is we only explore an approach for unlocking the inherent CL potential of LMs through updating the gradient by the magnitude of output. There exists more discussions on exploiting innate features such as activation sparsity as discussed in the Related Work section. Though we include our preliminary results in Table 11 of Appendix, the limitation can be further addressed in future work.
8 Acknowledgment
We thank all constructive comments from anonymous reviewers. Reynold Cheng and Wenyu Du were supported by the Hong Kong Jockey Club Charities Trust (Project 260920140), the University of Hong Kong (Project 109000579), the HKU Outstanding Research Student Supervisor Award 2022-23, and the HKU Faculty Exchange Award 2024 (Faculty of Engineering).
@@ B.1 Dataset Information
Method Standard CL Benchmark (4 tasks) Longer CL Benchmark (15 tasks)
@@
Table 8: Summary of the results on two standard CL benchmarks with T5-large model. Averaged accuracy after training on the last task is reported. All results in the last block are averaged over 3 runs. (We reuse the table template and experiment results from O-LoRA Wang et al. (2023b) to construct the results of the top two blocks, methods denoted with ). It is noticeable some baselines in some previous literature show significant lower performance than ours(SeqFT in block 1 v.s. FT in block 3), we assume this may due to different hyperparameter choice for baseline methods. To this end, we conducted a grid search that ranges in [1e-3, 3e-4, 1e-4, 3e-5] for FT baseline. We recorded our experiment results and provide in Table 9.
@@ B.2 Training orders
learning rate Order1 Order2 Order3 Avg
(Seq)FT in O-LORA 18.9 24.9 41.7 28.5
1e-3 23.0 21.7 21.8 22.2
3e-4 53.1 50.8 31.9 45.3
1e-4 74.4 75.0 77.5 75.7
3e-5 75.0 76.1 77.2 76.1
Table 9: The grid search for baseline FT. It is clearly that learning rates play an important role for CL.
@@ D.4 Visualization
To investigate how our method enhances model performance, we visualized the variation in product magnitudes between an FT model and an FT model augmented with our MIGU technique in Figures 24,32. We employed heatmaps to depict the similarity in product magnitude distributions across different tasks. Our findings reveal that task similarity in the FT model with MIGU implementation is markedly reduced. This suggests that the models trained with our method exhibit more distinctive weight activations for different tasks, thereby mitigating their conflict. This distinction in activation patterns indicates our method’s ability to foster more task-specific representations within the model, contributing to its improved performance across varied learning scenarios.
@@ E.2 Result & Analysis
Standard Benchmark Long Benchmark
FT 76.5 69.6
+ gradient-based 78.4 72.3
+ weight-based 77.2 71.3
+ MIGU 78.8 74.7
Table 11: Finding Important Weights Comparison
Appendix F More Ablation on Gradient Mask Threshold
We also used the original standard CL benchmark and plot all five curves of our approach (+MIGU) for gradient mask threshold from 0.0 to 0.9 in Section 4.1. The optimal threshold value for FT+MIGU, LoRA+MIGU, and IncLoRA+MIGU settings is 0.7 while LoRAReplay+MIGU is 0.4 as shown in Figure 19999The ablation on threshold search only reports one run, so it does not align with the results in Section 4.1.. OIncLoRA+MIGU is only 0.1, which may due to the parameter updating regularized by the OIncLoRA method itself. The optimal value for IncLoRA+MIGU is 0.6, close to FT+MIGU, LoRA+MIGU, and IncLoRA+MIGU settings. Surprisingly, with only 5% ($T=0.95$) or 1% ($T=0.99$) parameters updating, LoRA+MIGU still beats LoRA by a wide margin. This interesting finding may indicate that only a small proportion of proportional weights with large magnitudes is crucial for successful CL settings, which may be worth future investigation.